October 13, 2021 Calorimeter
Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time
- To join via a Web Browser, go to the page  https://bluejeans.com/907185247.
- To join via Polycom room system go to the IP Address: 184.108.40.206 (bjn.vc) and enter the meeting ID: 907185247.
- To join via phone, use one of the following numbers and the Conference ID: 907185247.
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
- More information on connecting to bluejeans is available.
Participant Direct Lines
- JLab Phone in CC L207 is 757-269-7084 (usual room)
- JLab Phone: in CC F326 is 757-269-6460
- Phone in the Regina Video-conference Suite is 306-585-4204
- Athens Phone: in Christina's office is 011-30-210-727-6947
- FCAL HDFCAL log book
- BCAL HDBCAL log book
- CCAL HDCCAL log book
- Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
- Photon Reconstruction Task Force
Goals for Calorimetry Group
- Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
- Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
- Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.
- Short term CALWG Items
- Long term items
- Calorimeter work packages
- AI - BCAL Monitoring:
- Run Operations
- Action Items
- Any other business
Attendees: Varun, Tolga, Mark, Sean, Elton, Rebecca, Jon, Susan, Matt, Malte, Igal, Zisis
- Announcements: DNP talks this week.
- Run Operations
- FCAL: Mark reported that basic maintenance continues. Earlier this morning, Igal showed his calibration dirk and the very good quality of the data. The hard work by IU+JLab has really paid off. Kudos to all participants.
- BCAL: is being used as a Veto for the PrimEx run.
- CCAL: no report.
- BCAL: Sean does the first monitoring pass (includes adjustments in timing) which will prepare files for Mark in a day or two. Mark will check attenuation lengths and a few other constants, and then Karthik does the pi0 gain calibration.
- BCAL (online): Varun gave un updated on the status of monitoring. Yost LEDs and channels are stable. There are still missing DN LED pulses and some UP. Some skim files were missing and Varun notified David, who restarted HOSS. Varun's study:" 39 DN and 3 UP LEDS give no response. 96 DN and 16 UP fluctuate (deviation greater than 5%). Varun will cross reference these LEDs with Sarah's report from 2020 data and will draw their locations so we can look for patterns in the BCAL.
- BCAL (offline): Tolga looked at Batch 1, Occupancy and Recon 1 (other plots are not present because of PrimEx running conditions for the BCAL). Sean commented that the first production run is 81267, so before that one should not look too closely. However, the ROOTSpy Reference plots need to be updated for shift takers. Marianna should be kept in the loop.
- FCAL (offline): Susan is at JLab and will go to the Counting House this afternoon (paperwork and badge are done). Susan is also looking at the PS and CCAL, in addition to FCAL.
- Reconstruction: no report.
- Simulations: no report.
- Action Items: Zisis updated these. No further changes needed at present.
- Any other business: none.